### v2

# Multi-Objective Coverage Bayesian Optimization (MOCOBO)

Abstract
In multi-objective black-box optimization, the goal is typically to find solutions that optimize a set of $T$ black-box objective functions, $f_{1},\ldots f_{T}$, simultaneously. Traditional approaches often seek a single Pareto-optimal set that balances trade-offs among all objectives. In this work, we consider a problem setting that departs from this paradigm: finding a small set of $K<T$ solutions, that collectively “covers" the $T$ objectives. A set of solutions is defined as “covering" if, for each objective $f_{1},\ldots f_{T}$, there is at least one good solution. A motivating example for this problem setting occurs in drug design. For example, we may have $T$ pathogens and aim to identify a set of $K<T$ antibiotics such that at least one antibiotic can be used to treat each pathogen. To address this problem, we propose Multi-Objective Coverage Bayesian Optimization (MOCOBO), a principled algorithm designed to efficiently find a covering set. We validate our approach through experiments on challenging high-dimensional tasks, including applications in peptide and molecular design, where MOCOBO is shown to find high-performing covering sets of solutions. The results show that the coverage of the $K<T$ solutions found by MOCOBO matches or nearly matches the coverage of $T$ solutions obtained by optimizing each objective individually. Furthermore, in in vitro experiments, the peptides found by MOCOBO exhibited high potency against drug-resistant pathogens, further demonstrating the potential of MOCOBO for drug discovery.
1 Introduction
Figure 1: Traditional multi-objective optimization for $T$ objectives might select any point along the Pareto frontier, but in some situations like this any Pareto optimal point performs poorly on at least one objective. In situations where multiple $K<T$ solutions are allowed ($\blacksquare$), we can sometimes optimize all objectives well.
Bayesian optimization (BO; 31, 54, 21) is a general framework for sample-efficient optimiziation of black-box functions. By using a probabilistic surrogate model, such as a Gaussian process, Bayesian optimization balances exploration and exploitation to identify high-performing solutions with a limited number of function evaluations. BO has been successfully applied in a wide range of domains, including hyperparameter tuning [56, 64], A/B testing [37], chemical engineering [26], drug discovery [47], and more.
In multi-objective optimization the typical assumption is that objectives inherently trade off with one another, motivating the search for Pareto-optimal solutions that balance these trade-offs. Multi-objective Bayesian optimization methods focus on identifying diverse (in objective value) sets of solutions along the Pareto front to capture these trade-offs [27, 3, 63, 33, 11, 57, 4]. Indeed, without such trade-offs, optimization would trivialize into simply optimizing all of the objectives.
However, in many real-world scenarios, the trade-offs between objectives can become extreme, with some objectives being nearly mutually exclusive. This makes it virtually impossible to achieve satisfactory performance across all objectives simultaneously. Figure 1 provides a $1D$ example of this demonstrating how the Pareto-optimal points may fail to achieve acceptable performance on all objectives. Such scenarios motivate an interesting generalization of the multi-objective setting that arises in situations where we can propose a small set of solutions that together optimize the multiple objectives and avoid the implicit constraints imposed by extreme trade-offs.
One compelling example arises in drug design. In this domain, the goal may be to develop antibiotics that treat a broad spectrum of pathogens–here, activity against the individual pathogens serve as the objectives. While it is very true that we seek to optimize the objectives simultaneously in the sense that we prefer to have broad-spectrum antibiotics, it is at the same time perfectly reasonable for one drug in the set to be ineffective against a few of the pathogens if it allows for greater potency against others. This flexibility enables a more targeted and efficient approach to therapeutic design, where our realistic goal is to have a small set of drugs that are together extremely lethal for the set of pathogens, without requiring each individual drug to be.
To formalize this, we consider the “coverage optimization" problem: the task of covering a set of $T$ objectives with $K$ solutions, where coverage means that for each objective $f_{1},\dots,f_{T}$, there exists at least one high-performing solution among the $K$. In the context of drug design, this translates to identifying $K$ drugs such that each of the $T$ pathogens is effectively addressed by at least one drug in the set. This problem setting diverges from traditional multi-objective optimization by prioritizing collective coverage over balancing trade-offs within individual solutions.
To address the coverage problem setting, we introduce and experimentally validate Multi-Objective Coverage Bayesian Optimization (MOCOBO), a new algorithm designed to solve the “coverage optimization" problem within the Bayesian optimization framework. MOCOBO efficiently identifies a small set of $K$ solutions that collectively cover $T$ objectives. Unlike traditional multi-objective Bayesian optimization methods that seek to balance trade-offs across all objectives for every solution, MOCOBO focuses on identifying specialized sets of solutions to meet practical coverage requirements.
Contributions
  1. 1.
    We consider the problem setting of finding a “covering set" of $K<T$ high-performing solutions such that each relevant objective $f_{t}\in\left\{f_{1},\ldots f_{T}\right\}$ is optimized by at least one solution. As far as we are aware, this is the first work in the Bayesian optimization (BO) literature to consider the coverage optimization problem setting.
  2. 2.
    We propose a local BO algorithm, MOCOBO, for this problem setting, that extends to high dimensional objectives, structured input domains, and large numbers of function evaluations.
  3. 3.
    We experimentally validate MOCOBO on challenging, high-dimensional optimization tasks including structured drug discovery over molecules and peptides, demonstrating its ability to consistently outperform state-of-the-art BO methods in identifying high-performing covering sets of solutions.
  4. 4.
    We demonstrate empirically that the small sets of $K<T$ solutions found by MOCOBO consistently nearly match the performance of a larger set of $T$ solutions individually optimized for each task.
  5. 5.
    We demonstrate the potential of MOCOBO for drug discovery using an in vitro experiment showing it produces potent antimicrobial peptides that cover 9 out of 11 drug resistant or otherwise challenging to kill pathogens, with moderate activity on 1 out of 11.
2 Background and related work
Bayesian optimization (BO).
BO is a method for sample-efficient black-box optimization. It operates iteratively, where each cycle involves training a probabilistic surrogate model, typically a Gaussian process (GP; 50), on data collected from the black-box objective function. BO then employs the predictive posterior $p(y\mid\mathbf{x},D)$ of the surrogate model to compute an acquisition function that gives a policy for choosing the next candidate(s) to evaluate, balancing exploration and exploitation.
Multi-Objective Bayesian optimization (MOBO).
Recent advances in MOBO have introduced numerous methods aimed at identifying diverse solutions along the Pareto front to balance trade-offs among multiple objectives [27, 3, 63, 33, 11, 57, 4]. These approaches, however, are not well-suited to “coverage optimization”, where the goal is to find a small set of $K<T$ solutions that collectively address all $T$ objectives. Despite this, we compare our proposed method to MORBO, developed by Daulton et al. [11], which is regarded as a state-of-the-art method for high-dimensional MOBO tasks.
Bayesian optimization over structured search spaces.
BO has recently been used for optimizing structured search spaces like molecules or amino acid sequences by employing latent space Bayesian optimization, which utilizes a variational autoencoder (VAE; 32, 52) to convert structured inputs into a continuous latent space for BO [23, 15, 61, 24, 55, 30, 57]. The VAE encoder, $\Phi(\mathbf{x})$, transforms structured inputs $\mathbf{x}$ into continuous latent vectors $\mathbf{z}$, enabling BO to operate in the latent space. Candidate latent vectors are decoded using the VAE decoder, $\Gamma(\mathbf{z})$, to generate structured outputs. Maus et al. [44] extended TuRBO [17] for latent space optimization by jointly training the surrogate model and VAE using variational inference.
Trust Region Bayesian Optimization (TuRBO).
Local Bayesian optimization [17], including TuRBO-$M$, is one recent successful approach to high-dimensional BO. TuRBO-$M$ uses $M$ parallel local optimization runs, each maintaining its own dataset $\mathcal{D}_{i}$ and surrogate model. Each local optimizer proposes candidates only within a hyper-rectangular trust region $\mathcal{T}_{i}$. The trust region $\mathcal{T}_{i}$ is a rectangular area of the input space $\mathcal{X}$, centered on the incumbent $\mathbf{x}^{+}_{i}$. Its side length $\ell_{i}$ is bound by $[\ell_{min},\ell_{max}]$. If an optimizer improves the incumbent for $\rho_{succ}$ consecutive iterations, $\ell_{i}$ expands to $\min(2\ell_{i},\ell_{max})$. If not improved in $\rho_{fail}$ iterations, $\ell_{i}$ is halved. Optimizers are restarted if $\ell_{i}$ drops below $\ell_{min}$. While TuRBO-$M$ is not directly applicable to our problem, we follow methods like MORBO [11] and ROBOT [45] in adapting the use of $M$ coordinated trust regions ($K$ in our case) to a different problem setting.
Prior work on coverage optimization and related problem settings.
To our knowledge, the coverage optimization problem setting considered in this paper has not been previously explored in the Bayesian optimization literature. However, similar problem settings have been studied in the context of gradient-based multi-objective optimization. In particular, Ding et al. [13], Li et al. [39], and Lin et al. [40] have investigated related problems in this domain. Ding et al. [13] demonstrate that the coverage optimization problem generalizes various clustering methods, including k-means clustering, and propose an algorithm that extends k-means++ [1] and Lloyd’s algorithm [43] to efficiently solve the coverage problem using gradient descent. Li et al. [39] introduce Many-objective Multi-solution Transport (MosT), a novel framework that scales multi-objective gradient-based optimization to a large number of objectives. MosT uses an optimal transport to establish a balanced assignment between objectives and solutions, allowing each solution to focus on a specific subset of objectives and ensuring that the collective set of solutions covers all objectives. Lin et al. [40] introduce a novel Tchebycheff set (TCH-Set) scalarization approach for covering multiple conflicting objectives using a small set of collaborative solutions. They further propose a smooth Tchebycheff set (STCH-Set) scalarization method to handle non-smoothness in TCH-Set scalarization. While these three papers address related coverage optimization problems, their proposed methods rely on differentiable objective functions, making them unsuitable for the black-box optimization problems considered in this paper. Liu et al. [42] were the first to consider the coverage problem in the setting of black-box optimization. They propose a clustering-based swarm optimization algorithm (CluSO) for the coverage problem in the setting of multi-objective black-box optimization. In Section 4, we provide empirical results demonstrating that our Bayesian optimization approach (MOCOBO) consistently outperforms CluSO by a large margin across all tasks considered.
3 Methods
We consider the task of finding a set of $K<T$ solutions $S^{*}=\left\{\mathbf{x}^{*}_{1},\ldots,\mathbf{x}^{*}_{K}\right\}$ that “covers" $T$ objectives $f_{1},\ldots,f_{T}$. Here, the set $S^{*}$ “covers" the $T$ objectives if, for each objective $f_{t}=f_{1},\ldots,f_{T}$, there is at least one solution $\mathbf{x}^{*}_{k}\in S^{*}$ for which $f_{t}$ is well optimized by $\mathbf{x}^{*}_{k}$. We evaluate how well a set of $K$ points $\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{K}\right\}$ “covers" the $T$ objectives using the “coverage score" $c$:
$c(\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{K}\right\})$ $=\sum_{t=1}^{T}\max_{k=1}^{K}f_{t}(\mathbf{x}_{k})$ (1)
Formally, we seek a set $S^{*}:=\left\{\mathbf{x}^{*}_{1},\ldots,\mathbf{x}^{*}_{K}\right\}$ such that:
$S^{*}$ $=\operatorname*{arg\,max}_{\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{K}\right\}\subseteq\mathcal{X}}c(\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{K}\right\})$ (2)
The coverage score in (1) only credits a single solution for each objective. If, for example, $f_{t}(\cdot)$ is maximized by one of the $\mathbf{x}_{i}$ in the solution set, improving the value of $f_{t}(\cdot)$ on some other $\mathbf{x}_{j\neq i}$ becomes irrelevant so long as $f_{t}(\mathbf{x}_{i})\geq f_{t}(\mathbf{x}_{j})$. In the setting where $K=1$, the coverage score collapses into a trivial linearization of the objectives and true multi-objective BO methods should be preferred. In the setting where $K=T$, the coverage score is trivially optimized by maximizing each objective independently.
3.1 Multi-Objective Coverage Bayesian Optimization (MOCOBO)
In this section, we propose MOCOBO - an algorithm which extends Bayesian optimization to the problem setting above. On each step of optimization $s$, we use our current set of all data evaluated so far $D_{s}=\left\{(\mathbf{x}_{1},\mathbf{y}_{1}),\ldots,(\mathbf{x}_{n},\mathbf{y}_{n})\right\}$ to define the best covering set $S^{*}_{D_{s}}=\left\{\mathbf{x}^{*(s)}_{1},\ldots,\mathbf{x}^{*(s)}_{K}\in D_{s}\right\}$ found so far. Here $\mathbf{y}_{i}=(f_{1}(\mathbf{x}_{i}),\ldots,f_{T}(\mathbf{x}_{i}))$ and $n$ is the number of data points evaluated so far at step $s$. Following Equation 2, we define $S^{*}_{D_{s}}$ as follows:
$S^{*}_{D_{s}}$ $=\operatorname*{arg\,max}_{\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{K}\right\}\subseteq D_{s}}c(\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{K}\right\})$ $=\operatorname*{arg\,max}_{\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{K}\right\}\subseteq D_{s}}\sum_{t=1}^{T}\max_{k=1}^{K}f_{t}(\mathbf{x}_{k})$ (3)
$c(S^{*}_{D_{s}})$ denotes the best coverage score found by the optimizer after optimization step $s$.
3.1.1 Candidate selection with Expected Coverage Improvement (ECI)
Given $S^{*}_{D_{s}}$, our surrogate model’s predictive posterior $p(y\mid\mathbf{x},D)$ induces a posterior belief about the improvement in coverage score achievable by choosing to evaluate at $\mathbf{x}$ next. We extend the typical expected improvement (EI) acquisiton function in the natural way to the coverage optimization setting by defining expected coverage improvement (ECI):
$\mathrm{ECI}(\mathbf{x})$ $=\mathbb{E}_{p(y|\mathbf{x},D)}[\max(0,c(S^{*}_{D_{s}\cup\left\{(\mathbf{x},\mathbf{y})\right\}})-c(S^{*}_{D_{s}}))]$ (4)
Here $c(S^{*}_{D_{s}})$ is the coverage score of the best possible covering set from among all data observed $D_{s}$ – as we shall see, constructing this set $S^{*}_{D_{s}}$ will be our primary challenge. $c(S^{*}_{D_{s}\cup\left\{(\mathbf{x},\mathbf{y})\right\}})$ is the coverage score of the best possible covering set after adding the observation $(\mathbf{x},\mathbf{y})$. Thus, ECI gives the expected improvement in the coverage score after making an observation at point $\mathbf{x}$. We aim to select points during acquisition that maximize ECI.
We estimate ECI using a Monte Carlo (MC) approximation. To select a single candidate $\hat{\mathbf{x}}$, we sample $m$ points $P=\left\{\mathbf{p}_{1},\mathbf{p}_{2},...,\mathbf{p}_{m}\right\}$. For each sampled point $\mathbf{p}_{j}$, we sample a realization $\hat{\mathbf{y}}_{j}=(\hat{f}_{1}(\mathbf{p}_{j}),\ldots,\hat{f}_{T}(\mathbf{p}_{j}))$ from the GP surrogate model posterior. We leverage these samples to compute an MC approximation to the ECI of each $\mathbf{p}_{j}$:
$\mathrm{CI}(\mathbf{p}_{j})$ $=\max(0,c(S^{*}_{D_{s}\cup\left\{(\mathbf{p}_{j},\hat{\mathbf{y}}_{j})\right\}}-c(S^{*}_{D_{s}})))$ (5)
Here, $c(S^{*}_{D_{s}\cup\left\{(\mathbf{p}_{j},\hat{\mathbf{y}}_{j})\right\}})$ is the approximation of the coverage score of the new best covering set if we choose to evaluate candidate $\mathbf{p}_{j}$, assuming the candidate point will have the sampled objective values $\hat{\mathbf{y}}_{j}$. We select and evaluate the candidate $\mathbf{p}_{j}$ with largest expected coverage improvement.
3.1.2 Greedy approximation of $S^{*}_{D_{s}}$
The candidate acquisition method in Section 3.1.1 utilizes the best covering set of $K$ points among all data collected so far, $S^{*}_{D_{s}}$. On each step of optimization $s$, MOCOBO must therefore construct $S^{*}_{D_{s}}$ from all observed data $D_{s}$, as $S^{*}_{D_{s}}$ may change on each step of optimization after new data is evaluated and added to $D_{s}$.
Lemma 3.1 (NP-hardness of Optimal Covering Set).
Let $T,K$ be finite positive integers such that $K<T$. Let $f_{1},\ldots,f_{T}$ be real valued functions. Let $D_{s}=\left\{(\mathbf{x}_{1},\mathbf{y}_{1}),\ldots,(\mathbf{x}_{n},\mathbf{y}_{n})\right\}$ be a dataset of $n$ real valued data points such that for all $\mathbf{x}_{i}$, $\mathbf{y}_{i}=(f_{1}(\mathbf{x}_{i}),\ldots,f_{T}(\mathbf{x}_{i}))$. Let $S^{*}_{D_{s}}$ be the optimal covering set of size $K$ in $D_{s}$ as defined in Equation 3. Then, constructing $S^{*}_{D_{s}}$ is NP-hard.
Proof.
See full proof of Lemma 3.1 in Section A.9. ∎
Algorithm 1 Greedy $(1-\frac{1}{e})$-Approximation for Finding $S^{*}_{D_{s}}$ (Incremental Strategy)
0:  Dataset $D_{s}$, observed values $f_{t}(\mathbf{x})$ for all $t\in\{1,\ldots,T\}$ and $\mathbf{x}\in D_{s}$, set size $K$.
  Initialize $A\leftarrow\emptyset$. {Start with an empty set.}
  Compute the initial coverage score for $A$: $c(A)\leftarrow 0$.
  for $k=1$ to $K$ do
     Initialize $\mathbf{x}_{\text{best}}\leftarrow\text{None}$ and $\Delta_{\text{best}}\leftarrow-\infty$.
     for $\mathbf{x}\in D_{s}\setminus A$ do
        Compute marginal coverage of adding $\mathbf{x}$ to $A$:
$\Delta c\leftarrow\sum_{t=1}^{T}\max\left(\max_{\mathbf{x}^{\prime}\in A}f_{t}(\mathbf{x}^{\prime}),f_{t}(\mathbf{x})\right)-c(A).$
        if $\Delta c>\Delta_{\text{best}}$ then
           $\mathbf{x}_{\text{best}}\leftarrow\mathbf{x}$ and $\Delta_{\text{best}}\leftarrow\Delta c$.
        end if
     end for
     Update $A\leftarrow A\cup\{\mathbf{x}_{\text{best}}\}$, $c(A)\leftarrow c(A)+\Delta_{\text{best}}$.
  end for
  Output: $A^{*}_{D_{s}}\leftarrow A$.
Since constructing $S^{*}_{D_{s}}$ is NP-hard, we use an approximate construction of $S^{*}_{D_{s}}$ on each step of optimization $s$. We present Algorithm 1 with $O(nKT)$ running time that provides a $(1-\frac{1}{e})$-approximation of $S^{*}_{D_{s}}$.
Theorem 3.2.
The set $A^{*}_{D_{s}}$ output by Algorithm 1 satisfies $c(A^{*}_{D_{s}})\geq\left(1-\frac{1}{e}\right)c(S^{*}_{D_{s}})$.
Proof.
See full proof of Theorem 3.2 in Section A.10. ∎
Algorithm 1 runtime.
The algorithm iterates $K$ times to construct the covering set. In each iteration, it evaluates at most $n$ candidate points, and for each candidate, it computes the incremental coverage score by iterating over $T$ objectives. The runtime is thus $O(K\cdot n\cdot T)$. For practical applications where we can assume relatively small $K$ and $T$, the runtime is approximately $O(n)$. For an empirical evaluation of the wall-clock runtime, see Section A.2.3.
Corollary 3.3 (Algorithm 1 is the best possible approximation of $S^{*}_{D_{s}}$).
There is no polynomial runtime algorithm that provides a better approximation ratio unless $P=NP$.
Proof.
This follows from the fact that we reduced from Max $k$-cover, for which a better approximation ratio is not practically achievable unless $P=NP$ [19]. ∎
3.1.3 Extending ECI to the batch acquisition setting (q-ECI)
In batch acquisition, we select a batch of $q>1$ candidates for evaluation. Following recent work on batch EI [71, 67], we define q-ECI, a natural extension of ECI to the batch setting:
$q\text{-ECI}(\mathbf{X})$ $=\mathbb{E}_{p(\mathbf{Y}|\mathbf{X},D)}\left[\max(0,\max_{r=1,\dots,q}c\left(S^{*}_{D_{s}\cup\left\{(\mathbf{x}_{r},\mathbf{y}_{r})\right\}}\right)-c\left(S^{*}_{D_{s}})\right)\right]$ (6)
q-ECI gives the expected improvement in coverage score after simultaneously observing the batch of $q$ points $\mathbf{X}=\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{q}\right\}$. However, the resulting Monte Carlo expectation would require $O(q\times m)$ evaluations of Algorithm 1. We therefore adopt a more approximate batching strategy for practical use with large $q$ (see Section A.3 for details).
3.1.4 MOCOBO trust regions
In order to find a set of $K$ solutions, MOCOBO maintains $K$ simultaneous local optimization runs using $K$ individual trust regions. Each local run $k\in\left\{1,\ldots,K\right\}$ aims to find a single solution $\mathbf{x}^{*}_{k}$, which together form the desired set $S^{*}$. As in the original TuRBO paper, trust regions are rectangular regions of the search space $\mathcal{T}_{k}\subseteq\mathcal{X}$ defined solely by their size and center point. We center the $K$ trust regions on the best covering set of solutions observed so far during optimization. In particular, on each optimization step $s$, we construct an approximation of $S^{*}_{D_{s}}$ using Algorithm 1, and center each trust region $\mathcal{T}_{k}$ on the corresponding point $\mathbf{x}^{*(s)}_{k}$ in $S^{*}_{D_{s}}=\left\{\mathbf{x}^{*(s)}_{1},\ldots,\mathbf{x}^{*(s)}_{K}\right\}$. We then use our proposed ECI (or q-ECI when $q>1$) acquisition function to select $q$ candidates for evaluation from within each trust region. The MC approximation of ECI described in Section 3.1.1 can be straightforwardly applied to select candidates in trust region $\mathcal{T}_{k}$ by sampling the $m$ discrete points from within the rectangular bounds of $\mathcal{T}_{k}$. Since we select and evaluate $q$ candidates from each of the $K$ trust regions, the total number of observed data points $n$ at step $s$ of MOCOBO is $n=s\times K\times q$. As in the original TuRBO algorithm, each trust region $\mathcal{T}_{k}$ has success and failure counters that dictate the size of the trust region. For MOCOBO, we count a success for trust region $\mathcal{T}_{k}$ whenever $\mathcal{T}_{k}$ proposes a candidate on step $s$ that improves upon the best coverage score and is included in $S^{*}_{D_{s+1}}$.
4 Experiments
Figure 2: Coverage optimization results on all tasks considered.
We evaluate MOCOBO on four high-dimensional, multi-objective BO tasks for which finding a set of $K<T$ solutions to cover the $T$ objectives is desirable. Detailed descriptions of each task are in Section 4.1. Two tasks involve continuous search spaces, allowing direct application of MOCOBO, while the other two involve structured spaces (molecules and peptides), requiring an extension for structured optimization.
Implementation details and hyperparameters.
We implement MOCOBO using BoTorch [2] and GPyTorch [20]. Code to reproduce all results in the paper is available on GitHub: placeholder. We use an acquisition batch size of $20$ for all tasks and across all BO methods compared. Further implementation details are provided in subsection A.7. For our structured optimization tasks, we apply MOCOBO in the latent space of a pre-trained VAE on which we perform regular end-to-end updates with the surrogate model during optimization [44].
Plots.
In Figure 2, we plot the best coverage score $c$ Equation 1 obtained by the $K$ best covering solutions found so far after a certain number of function evaluations on each task. Since BO baseline methods are not designed to optimize coverage directly, instead aiming to find a single solution or a set of solutions to optimize the $T$ objectives, we plot the best coverage score $c$ obtained by the $K$ best covering solutions found by the method. All plots show mean coverage scores averaged over $20$ replications of each method, and show standard errors.
Baselines.
In all plots, we compare MOCOBO against CluSO, TuRBO, ROBOT, and MORBO [42, 17, 45, 11]. We also compare to sampling uniformly at random in the search space. For each method, we compare to the performance of the best set of $K$ solutions found.
Extending baselines to the structured BO setting.
In the case of the two structured optimization tasks (molecule and peptide design), we apply LOL-BO and ROBOT as described by design for these problem settings. For MORBO and CluSO, we take the straightforward approach of applying the methods directly in the continuous latent space of the VAE model without other adaptations. The same pre-trained VAE model is used across methods compared.
Peptide Amino Acid Sequence B1(-) B2(-) B3(-) B4(-) B5(-) B6(-) B7(-) B8(+) B9(+) B10(+) B11(+)
KKKKLKLKKLKKLLKLLKRL 1.017 1.040 1.893 0.999 8.613 0.966 1.039 65.999 38.361 338.692 1.393
IFHLKILIKILRLL 0.999 15.565 1.860 1.952 404.254 486.860 406.034 1.233 1.318 7.359 0.981
SKKIKLLGLALKLLKLKLKL 2.654 3.268 3.113 4.854 4.923 12.967 14.610 22.631 29.685 254.306 3.947
KKKKLKLKKLKRLLKLKLRL 0.939 0.906 1.124 1.310 10.909 1.384 1.711 12.776 32.884 434.193 1.037
Table 1: Best set of $K=4$ peptides found by one run of MOCOBO for the “template free" peptide design task. We provide the APEX model’s predicted MIC for each sequence on each of the $11$ target pathogenic bacteria. The target pathogenic bacteria B1,,$\ldots$,, … , B11 are listed in Table A.7. (-) and (+) indicate Gram negative and Gram positive pathogenic bacteria respectively. The best/lowest MIC achieved for each pathogenic bacteria is in bold in each column. See row “TF1" in Figure 3 left, and Figure A.1 for in vitro MICs for this set of $K=4$ peptides. See Table A.1 for an analogous result on the “template constrained" peptide design task.
Extending baselines to the coverage optimization setting.
As TuRBO, LOL-BO, and ROBOT target single-objective optimization, we conduct $T$ independent runs to optimize each of the $T$ objectives for the multi-objective task. We use all solutions gathered from the $T$ runs to compute the best covering set of $K$ solutions found by each method. Unlike LOL-BO and TuRBO which seek a single best solution for a given objective, a single run of ROBOT seeks a set of $M$ solutions that are pairwise diverse. We run ROBOT with $M=K$ so that each independent run of ROBOT seeks $M=K$ diverse solutions. The aggregate result of the $T$ independent runs for ROBOT is thus $K$ diverse solutions for each of the $T$ objectives. We compare to the best covering $K$ solutions from among those $K*T$ solutions. See Section A.7.4 for more details on the diversity constraints used by ROBOT and the associated hyperparameters. MORBO is a multi-objective optimization method and can thus be applied directly to each multi-objective optimization task. For each run of MORBO, we compare to the best covering set of $K$ solutions found among all solutions proposed by the run. For CluSO, no extension is needed as this method is designed for coverage optimization. We run CluSO with the same $K$ used by MOCOBO, and all other hyperparameters set to CluSO defaults.
T individually optimized solutions baseline.
We also compare to a brute-force method involving $T$ separate single-objective optimizations for each of the $T$ objectives, using TuRBO for each run or LOL-BO for molecule and peptide design. This is not an alternative for finding $K<T$ covering solutions, but instead identifies $T$ solutions, one per objective, approximating a ceiling we can achieve on performance without the limit of $K<T$ solutions. Approaching the performance of this baseline implies that we can find $K$ solutions that do nearly as well as if we were allowed $T$ solutions instead.
4.1 Tasks
Peptide design.
[…]
Template free vs template constrained peptide design.
We evaluate MOCOBO on two variations of the peptide design task: “template free" (TF) and “template constrained" (TC). For TF, we allow the optimizer to propose any sequence of amino acids. For TC, we add a constraint that any sequence proposed by the optimizer must have a minimum of $75$ percent sequence similarity to at least one of the $10$ template amino acid sequences in Table A.8. These $10$ templates were mined from extinct organisms and selected by Wan et al. [66]. The motivation of the template constrained task is to design peptides specifically likely to evade antibiotic resistance by producing “extinct-like” peptides that bacteria have not encountered in nature in thousands of years. For MOCOBO and all BO baselines (TuRBO, LOL-BO, ROBOT, and MORBO), we handle the optimization constraint by adapting techniques from SCBO [16]. For CluSO and random sampling, we handle the constraint with rejection sampling.
Ranolazine MPO molecule design.
Ranolazine is a drug used to treat chest pain. The original Ranolazine MPO task from the Guacamol benchmark suite of molecular design tasks [6] aims to design an alternative to this drug: a molecule with a high fingerprint similarity to Ranolaize that includes fluorine. We extend this task to the multi-objective optimization setting by searching for alternatives to Ranolazine that include $T=6$ reactive nonmetal elements not found in Ranolazine: fluorine, chlorine, bromine, selenium, sulfur, and phosphorus. We aim to cover the $T=6$ objectives with $K=3$ molecules. We use the SELFIES-VAE introduced by Maus et al. [44] to map the molecular space to a continuous $256$ dimensional space.
Figure 3: (Left) In vitro results for the two best “template free" (TF1, TF2) and two best “template constrained" (TC1, TC2) runs of MOCOBO for the peptide design task. Columns are the best/lowest in vitro MIC among the $K=4$ peptides found by MOCOBO for each target pathogenic bacteria B1,,$\ldots$,, … , B11 listed in Table A.7. (-) and (+) indicate Gram negative and Gram positive respectively. TF1 and TC1 correspond to the single runs of MOCOBO shown in Table 1 and Table A.1 respectively. Methods used to obtain in vitro MICs are provided in Section A.1. (Right) Ablation study comparing MOCOBO to optimization performance where a known “good” partitioning of the $T$ objectives into $K$ subsets is available in advance. We individually optimize $K$ solutions, one for each partition.
Rover.
The rover optimization task introduced by Wang et al. [68] consists of finding a $60$-dimensional policy that allows a rover to move along some trajectory while avoiding a fixed set of obstacles. To frame this as a multi-objective optimization task, we design $T$ unique obstacle courses for the rover to navigate. The obstacle courses are designed such that no single policy can successfully navigate all courses. We seek $K<T$ policies so that at least one policy enables the rover to avoid obstacles in each course. We evaluate on three unique instances with varying numbers of obstacle courses. For the instances of this task with $T=4$ and $T=8$, we seek to cover the objectives with $K=2$ solutions. For the $T=12$ instance, we seek to cover the objectives with $K=4$ solutions.
Image tone mapping.
In high dynamic range (HDR) images, some pixels (often associated with light sources) can dominate overall contrast, requiring adjustments to reveal detail in low-contrast areas, a problem known as tone mapping [51, Section 6]. Tone mapping algorithms involve various tunable parameters, resulting in a high-dimensional optimization problem of subjectively perceived quality. We seek a covering set of $K=4$ solutions to optimize a set of $T=7$ image aesthetic (IAA) and quality (IQA) assessment metrics from the pyiqa library [9] (see metrics listed in Table A.5). Our practical goal is that, while we do not know a priori which metric is best for a particular image, covering all metrics may result in at least one high quality image. We optimize over the $13$-dimensional parameter space of an established tone mapping pipeline to tone-map the “Stanford Memorial Church" [12] and “desk lamp" [7] benchmark images. See Section A.8.1 for details.
4.2 Optimization results
In Figure 2, we provide optimization results comparing MOCOBO to the baselines discussed above on all tasks. The results show that MOCOBO finds sets of $K$ solutions that achieve higher coverage scores across tasks. The “T Individually Optimized Solutions" baseline appears as a horizontal dotted line in all plots of Figure 2, representing the average coverage score of $T$ individually optimized solutions, serving as an approximation of the best possible performance without the constraint of a limited $K<T$ solution set. Results in Figure 2 demonstrate that the smaller set of $K<T$ solutions identified by MOCOBO nearly equals the performance of the complete set of $T$ individually optimized solutions. This result depends on using domain knowledge to choose $K$ large enough to achieve it. Results in Figure 2 show that matching the performance of $T$ optimized solutions is possible with some values of $K\ll T$.
Peptide design results.
In Table 1, we provide the $K=4$ peptides found by one run of MOCOBO for the template free (TF) peptide design task. For each peptide, we provide the the APEX 1.1 model’s predicted MIC for each of the $11$ target bacteria. MIC values $\leq 16$ $\mu$ mol $L^{-1}$ are considered to be “highly active" against the target pathogenic bacteria [66]. We highlight in Table 1 the comparison between the second peptide with the other three peptides. B1-B7 are Gram negative (GN) bacteria, while B8-B11 are Gram positive (GP). Peptide 2 specialized to the GP bacteria (B8-B11 scores predicted highly active compared to the other peptides) at the expense of broad spectrum activity for the GN bacteria (B5, B6, B7 predicted inactive). By specializing its solutions to GN and GP bacteria separately, MOCOBO achieves low MIC across the target bacteria with only $K=4$ peptides. A similar table of results is also provided for the template constrained (TC) peptide design task (see Table A.1).
Figure 4: (Four Leftmost Panels) $T=4$ varied obstacle courses as multiple objectives for the rover task and $K=2$ covering trajectories (blue, magenta) found by MOCOBO. The first trajectory (magenta) successfully navigates obstacle courses 2 and 3 (second and third panels from the left). The second trajectory (blue) navigates obstacle courses 1 and 4 (first and fourth panels from the left). (Middle Panel) Initial hdr desk image for the desk variation of the “image tone mapping" task. (Four Rightmost Panels) Images obtained by transforming the hdr desk image using the best covering set of $K=4$ solutions found by a single run of MOCOBO.
In Figure 3 left, we provide in vitro results for the two best TF and two best TC runs of MOCOBO for the peptide design task. Here, “best" means runs that achieved highest coverage scores according to the APEX 1.1 model. For each run, Figure 3 left provides the best/lowest in vitro MIC among the $K=4$ peptides found by MOCOBO for each target bacteria. Results demonstrate that solutions found by MOCOBO optimizing against the in-silico APEX 1.1 model achieve good coverage of the $11$ target pathogenic bacteria in vitro. In Figure A.1, we provide the full set of in vitro results for each these runs of MOCOBO, with MIC values obtained by each of the $K=4$ peptides found for each target bacteria. Methods used to obtain in vitro MICs are provided in Section A.1.
Molecule design results.
In Table A.2, we provide results for the $K=3$ molecules found by one run of MOCOBO for the molecule design task. Each of the $6$ target elements is successfully present in one of the $K=3$ molecules in the best covering set. These three molecules therefore effectively cover the $T=6$ objectives, with each objective having a max score $>0.9$.
Rover results.
In Figure 4 left, we depict the $4$ varied obstacle courses used for the $T=4$ variation of the rover task and $K=2$ covering trajectories found by a single run of MOCOBO. In Figure A.3 and A.4, we provide analogous figures for the $T=8$ and $T=12$ variations. In each example, the MOCOBO optimized set of trajectories covers all obstacle courses such that all obstacles are avoided.
Image tone mapping results.
In Figure 4 right and Figure A.2, we provide the original hdr images and the images obtained by the $K=4$ solutions found by a single run of MOCOBO for the church and desk test images respectively. In both variations, three of the four solutions result in high quality images, while one results in a poor quality image (see middle right church image and rightmost desk image). The poor quality images were each generated by the one solution that MOCOBO used to cover metrics $5$ and $6$ (see metric ID numbers in Table A.5). This indicates that metrics $5$ and $6$ are poor indicators of true image quality in this setting. This result highlights a useful application of MOCOBO: By dedicating one of the $K$ solutions to maximizing the misleading metrics ($5$ and $6$), the other three solutions to can focus on the remaining subjectively better metrics.
4.3 Ablation study
A challenging aspect of our problem setting is that we do not know a priori the best way(s) to divide the $T$ objectives into $K$ subsets such that good coverage can be obtained. In this section, we seek to answer the question: what is the efficiency lost by MOCOBO due to not knowing an efficient partitioning of the objectives in advance?. To construct a proxy “efficient” partition to measure this, we first run MOCOBO to completion on a task, and then consider optimization as if we had known the partitioning of objectives found at the end in advance.
With such an “oracle” partitioning in hand ahead of time, we can efficiently optimize by running $K$ independent optimization runs: one to find a single solution for each subset in the fixed, given partition. With a fixed partition, the coverage score in (1) reduces to optimizing the sum of objectives in each relevant subset independently. In Figure 3 right, we compare this strategy directly to MOCOBO on three optimization tasks. We plot the best coverage score obtained by $K$ independent runs of TuRBO/LOL-BO on the fixed oracle $K$-partitioning of the objectives by combining their solutions into a single set. MOCOBO achieves the same average best coverage scores with minimal loss in optimization efficiency despite having to discover an efficient partitioning during optimization. See Section A.2.2 for an additional ablation study of MOCOBO when all objectives are highly conflicting.
5 Discussion
By bridging the gap between traditional multi-objective Bayesian optimization (BO) and practical coverage requirements, MOCOBO offers an effective approach to tackle critical problems in drug design and beyond. This framework extends the reach of Bayesian optimization into new domains, providing a robust solution to the challenges posed by extreme trade-offs and the need for specialized, collaborative solutions. See Section A.4 for additional discussion and limitations of MOCOBO.
Acknowledgments and Disclosure of Funding
N. Maus was supported by the National Science Foundation Graduate Research Fellowship; K. Kim was supported by a gift from AWS AI to Penn Engineering’s ASSET Center for Trustworthy AI; J. R. Gardner was supported by NSF awards IIS-2145644 and DBI-2400135; C. de la Fuente-Nunez was supported by NIH grant R35GM138201, and by Defense Threat Reduction Agency grants HDTRA11810041, HDTRA1-21-1-0014, and HDTRA1-23-1-0001.
Appendix A Appendix
A.1 Methods for obtaining in vitro minimal inhibitory concentration (MIC) data
In this section, we provide the methods used to produce all in vitro minimal inhibitory concentration (MIC) values reported in this paper.
Peptide synthesis.
All peptides were synthesized by solid-phase peptide synthesis using the Fmoc strategy and purchased from AAPPTec.
Bacterial strains and growth conditions used in the experiments.
[…]
Antibacterial assays.
The in vitro antimicrobial activity of the peptides was assessed by using the broth microdilution assay [8]. Minimal inhibitory concentration (MIC) values of the peptides were determined with an initial inoculum of $2\times 10^{6}$ cells $mL^{-1}$ in LB in microtiter $96$-well flat-bottom transparent plates. Aqueous solutions of the peptides were added to the plate at concentrations ranging from $0.0625$ to $64$ $\mu molL^{-1}$. The lowest concentration of peptide that inhibited 100 percent of the visible growth of bacteria was established as the MIC value in an experiment of $20h$ of exposure at 37 degrees Celsius. The optical density of the plates was measured at $600nm$ using a spectrophotometer. All assays were done as three biological replicates.
A.2 Additional results
In this section, we provide all additional empirical results not included in the main text.
A.2.1 Additional covering sets of solutions found by MOCOBO
In this section, we provide additional examples of covering sets of solutions found by MOCOBO for various tasks from Section 4.
[…]
In Figure A.1, we provide In vitro results for the two best “template free" and two best “template constrained" runs of MOCOBO for the peptide design task. Here, “best" means runs that achieved highest coverage scores according to the APEX 1.1 model. Figure A.1 provides in vitro MICs for each of the $K=4$ peptides found by each of these runs of MOCOBO, for each of the $11$ target pathogenic bacteria. Methods used to obtain in vitro MICs are provided in Section A.1.
Peptide Amino Acid Sequence B1(-) B2(-) B3(-) B4(-) B5(-) B6(-) B7(-) B8(+) B9(+) B10(+) B11(+)
KKLKIIRLLFK 18.594 17.067 4.278 5.352 13.460 50.442 24.543 456.831 431.276 441.292 20.305
WAIRGLKLATWLSLNNKF 6.771 20.358 14.644 10.477 65.172 97.404 59.195 19.846 33.459 237.697 7.708
RWARNLVRYVKWLKKLKKVI 2.171 4.589 2.641 3.073 54.400 11.444 19.150 75.588 89.977 413.386 2.913
HWITIAFFRLSISLKI 225.260 346.589 56.583 58.253 458.963 475.616 538.352 293.852 338.047 34.230 22.153
Table A.1: The best set of $K=4$ peptide sequences found by one run of MOCOBO for the “template free" peptide design task described in Section 4.1. For each of the four sequences, we provide the MIC according to the APEX model for each of the $11$ target pathogenic bacteria. The target pathogenic bacteria B1,,$\ldots$,, … , B11 are listed in Table A.7. (-) and (+) indicate Gram negative and Gram positive pathogenic bacteria respectively. The best/lowest MIC achieved for each pathogenic bacteria is in bold in each column. See row “TC1" in Figure 3 left, and Figure A.1 for in vitro MICs for this set of $K=4$ peptides.
In Table A.1, we provide an example of a covering set of peptides found by MOCOBO for the “template constrained" variation of the peptide design task.
In Figure A.2, we provide the original hdr image (leftmost panel), and the $K=4$ images produced using the $K=4$ solutions found by a single run of MOCOBO for the church image variation of the image tone mapping task. An analogous result for the desk image variation of the image tone mapping task can be found in the main text in Figure 4 right. A notable apparent limitation of the existing IAA/IQA metrics we used for the image tone mapping tasks is that they all favored monochromatic images for both the church and desk images. As such, the tone-mapped images obtained appear less colorful than their hand-tuned counterparts reported in prior work.
In Table A.2, we provide an example of a covering set of $K=3$ molecules found by MOCOBO for the molecule design task. As mentioned in Section 4, these three molecules effectively cover the $T=6$ objectives, as evidenced by the presence of all $T=6$ target elements in one of the $K=3$ molecules designed by MOCOBO.
In Figure A.3 and Figure A.4, we provide examples of a covering set of trajectories found for the $T=8$ and $T=12$ variations of the rover task respectively. In each figure, a panel is shown for each of the $T$ obstacles courses, with obstacles colored in red. The required starting point for the rover is a green point in the bottom left of each panel. The “goal" end point that the rover aims to reach without hitting any obstacles is the red point in the top right of each panel. Each panel also shows the best among the $K$ trajectories found by a run of MOCOBO for navigating each obstacle course. An analogous plot for the $T=4$ variation of the rover optimization task is provided in the main text in Figure 4.
Figure A.2: (Leftmost Panel) Image obtained by naively compressing the dynamic range of the HDR church image in the “image tone mapping" task. (Four Rightmost Panels) Images obtained by applying tone mapping to the church image using the best covering set of $K=4$ solutions found by a single run of MOCOBO.
Molecule (Smiles String) Obj 1 (add F) Obj 2 (add Cl) Obj 2 (add Br) Obj 2 (add Se) Obj 2 (add S) Obj 2 (add P)
CC=C(C)C(OC(=O)C(O)CCCCCCC(=O)O)
=CC=CCCCCCC[Se]CC(=O)NC1=CC=CC=C1C 0.8038 0.8038 0.8038 0.9108 0.8038 0.8038
CC=C(C)C(OC(=O)CCCCCCC(O)C(=S)Cl)
=CC=CCOCCCCC(O)CC(=O)NC1=CC=CC=C1C 0.8043 0.9114 0.8043 0.8043 0.9114 0.8043
CC=C(C)C(OC(=O)C(O)CCCCCCC(=O)CBr)
=CC=COCPCCCCN(C)CC(=O)[NH1]C1=CC=C(F)C=C1C 0.9097 0.8028 0.9097 0.8028 0.8028 0.9097
Table A.2: The best set of $K=3$ molecules found by one run of MOCOBO for the Ranolazine MPO multiple element molecule design task described in Section 4.1. For each of the three molecules, we provide the objective value obtained for each of the $T=6$ objectives. The best/highest objective value is in bold in each column. Each objective aims to add a different target element to Ranolazine (F, Cl, Br, Se, S, and P). The $T=6$ target elements are in bold in the smiles string [69] representation of each molecule.
Figure A.3: The 8 panels depict the 8 obstacle courses that the rover must navigate for the $T=8$ variation of the rover task, with obstacles colored in red. The required starting point for the rover is a green point in the bottom left of each panel. The “goal" end point that the rover aims to reach without hitting any obstacles is the red point in the top right of each panel. The line in each panel shows the best trajectory for navigating the obstacle course from among the $K=2$ covering trajectories found by a single run of MOCOBO. The first trajectory in the covering set is shown in magenta and successfully navigates obstacle courses 5, 6, 7, and 8 (Bottom Row). The second is shown in blue and successfully navigates obstacle courses 1, 2, 3, and 4 (Top Row).
Figure A.4: The 12 panels depict the 12 obstacle courses that the rover must navigate for the $T=12$ variation of the rover task, with obstacles colored in red. The required starting point for the rover is a green point in the bottom left of each panel. The “goal" end point that the rover aims to reach without hitting any obstacles is the red point in the top right of each panel. The line in each panel shows the best trajectory for navigating each obstacle course from among the $K=4$ covering trajectories found by a run of MOCOBO. The first trajectory in the covering set is shown in magenta and successfully navigates obstacle courses 6 and 7 (Middle Row: Center Left and Center Right Panels). The second is shown in blue and successfully navigates obstacle courses 2 and 8 (Top Row: Center Left Panel, Middle Row: Rightmost Panel). The third is shown in purple and successfully navigates obstacle courses 4, 5, and 11 (Top Row: Rightmost Panel, Middle Row: Leftmost Panel, and Bottom Row: Center Right Panel). The fourth trajectory is shown in orange and successfully navigates obstacle courses 1, 3, 9, 10, and 12 (Top Row: Leftmost Panel and Center Right Panel, Bottom Row: Leftmost, Center Left, and Rightmost Panels).
A.2.2 Additional ablation study: $T$ highly conflicting objectives
MOCOBO is designed for cases where some, but not all, of the objectives are highly conflicting, and this is the case for all of the tasks considered in Section 4. The conflicting pairs of the objectives prevent a single solution from optimizing all objectives well. The fact that not all pairs of objectives are completely conflicting, is the reason why it is possible to cover all T objectives with a small set of K solutions. If all pairs of the T objectives are highly conflicting, there is by definition no possible set of K<T solutions such that all objectives are well optimized. In this case, it’s best to use T individually optimized solutions if your goal is to find at least one solution that well optimizes each objective. However, it is still interesting to consider the performance of MOCOBO in the setting of T objectives that are pairwise highly conflicting. To investigate this, we construct a new variation of the multi-objective rover task described in Section 4.1. We design T=3 obstacle courses such that it is impossible for the rover to take any single path that avoids all obstacles in any pair of the obstacle courses. This results in T=3 objectives that are pairwise highly conflicting. We then run 20 replications of MOCOBO on this problem with K=2, asking MOCOBO to design K=2 solutions that cover the T=3 pairwise highly conflicting objectives. For this task, MOCOBO got an average coverage score of $2.932$, while T individually optimized solutions got an average coverage score of $12.899$. It is unsurprising that T individually optimized solutions achieved better coverage here since all three objectives are pairwise highly conflicting. However, this result demonstrates that MOCOBO is able to design sets of K=2 solutions that still achieve fairly high coverage scores.
Figure A.5: The three panels depict the T=3 obstacle courses that the rover must navigate for the ablation variation of the rover task where all T=3 objectives pairwise highly conflicting. Obstacles are colored in red. The three obstacle courses are designed such that any single path that successfully avoids all obstacles in one obstacle course, must hit some obstacles in both of the other obstacle courses. The required starting point for the rover is a green point in the bottom left of each panel. The “goal" end point that the rover aims to reach without hitting any obstacles is the red point in the top right of each panel. The line in each panel shows the best trajectory for navigating the obstacle course from among the K=2 covering trajectories found by a single run of MOCOBO. The first trajectory in the covering set is shown in magenta and is the best trajectory for navigating obstacle courses 2 and 3 (Middle Panel and Rightmost Panels). The second is shown in blue and successfully navigates obstacle course 1 (Leftmost Panel).
Figure A.5 provides a diagram of the T=3 pairwise highly conflicting obstacle courses, and an example of one of the covering sets of K=2 trajectories found by a single run of MOCOBO. One optimized trajectory (shown in blue) is designed by MOCOBO such that it specializes to navigate the first obstacle course, navigating it without hitting any obstacles (Leftmost Panel). The other optimized trajectory (shown in magenta) is designed by MOCOBO such that it avoids all obstacles in the third obstacle course (Rightmost Panel), while also minimizing total amount of impact with the obstacles in the second obstacle course (Middle Panel). Since it is by-design impossible to avoid all obstacles in all T=3 obstacle courses with only K=2 solutions, Figure A.5 demonstrates that MOCOBO was able to successfully balance trade-offs among the objectives, designing a set of K=2 solutions that minimized the total amount of obstacle impact across the three obstacle courses.
A.2.3 Wall-clock runtime of Algorithm 1
In Table A.3, we provide an empirical evaluation of the average wall-clock runtime of Algorithm 1 with different values of $N$ (the number of data points), $K$ (the covering set size), and $T$ (the number of objectives). We provide average run times with $N=20,200,2000,20000,200000,$ and $2000000$ for each combination of $K$ and $T$ that we ran MOCOBO on in Section 4. To improve the speed of Algorithm 1, in all experiments, we compute the inner for loop (lines 5-6 in Algorithm 1) in parallel. This involves simply computing the marginal coverage of adding all $N$ data points at once in parallel, and then updating $A$ with the point that achieved maximum marginal converge improvement. Parallel computation of the marginal coverage improvement of the $N$ points is straightforward to implement using PyTorch. In Table A.3, the “Runtime" column gives the average runtime of Algorithm 1 with this parallel computation of the of the inner for loop, and the “Runtime No Parallelization" column gives the average runtime without the parallel computation. Notice that as $N$ grows large, this parallel computation becomes essential to achieve the reasonably fast run times of Algorithm 1 needed to run MOCOBO. Average run times reported in Table A.3 are computed by running Algorithm 1 on a NVIDIA RTX A5000 GPU and averaging over $10$ runs.
N K T Runtime No Parallelization (seconds) Runtime (seconds)
20 2 4 $0.00605\pm 0.00385$ $0.00864\pm 0.00795$
200 2 4 $0.0230\pm 0.00391$ $0.00901\pm 0.00828$
2000 2 4 $0.194\pm 0.00389$ $0.00850\pm 0.00781$
20000 2 4 $2.0523\pm 0.0259$ $0.00852\pm 0.00780$
200000 2 4 $20.845\pm 0.0446$ $0.00872\pm 0.00796$
2000000 2 4 $195.740\pm 0.114$ $0.00923\pm 0.00792$
20 2 8 $0.00596\pm 0.00375$ $0.00906\pm 0.00833$
200 2 8 $0.0269\pm 0.00446$ $0.00894\pm 0.00808$
2000 2 8 $0.197\pm 0.00386$ $0.00865\pm 0.00794$
20000 2 8 $1.947\pm 0.00348$ $0.00864\pm 0.00789$
200000 2 8 $19.508\pm 0.00964$ $0.00915\pm 0.00825$
2000000 2 8 $193.317\pm 0.218$ $0.00962\pm 0.00785$
20 4 12 $0.00861\pm 0.00430$ $0.00873\pm 0.00785$
200 4 12 $0.0440\pm 0.00371$ $0.00875\pm 0.00789$
2000 4 12 $0.386\pm 0.00642$ $0.00910\pm 0.00821$
20000 4 12 $4.303\pm 0.120$ $0.00869\pm 0.00776$
200000 4 12 $38.596\pm 0.0895$ $0.00889\pm 0.00784$
2000000 4 12 $395.325\pm 0.169$ $0.0111\pm 0.00772$
20 3 6 $0.00704\pm 0.00384$ $0.00887\pm 0.00809$
200 3 6 $0.0334\pm 0.00372$ $0.00840\pm 0.00764$
2000 3 6 $0.314\pm 0.00654$ $0.00897\pm 0.00817$
20000 3 6 $2.942\pm 0.00478$ $0.00872\pm 0.00785$
200000 3 6 $28.732\pm 0.0156$ $0.00862\pm 0.00775$
2000000 3 6 $312.825\pm 0.203$ $0.00985\pm 0.00799$
20 4 7 $0.00765\pm 0.00370$ $0.00880\pm 0.00793$
200 4 7 $0.0448\pm 0.00365$ $0.00847\pm 0.00760$
2000 4 7 $0.412\pm 0.00252$ $0.00887\pm 0.00798$
20000 4 7 $3.917\pm 0.00461$ $0.00882\pm 0.00789$
200000 4 7 $38.932\pm 0.0402$ $0.00889\pm 0.00791$
2000000 4 7 $389.272\pm 0.588$ $0.0105\pm 0.00809$
20 4 11 $0.00772\pm 0.00377$ $0.00955\pm 0.00853$
200 4 11 $0.0422\pm 0.00383$ $0.00859\pm 0.00775$
2000 4 11 $0.387\pm 0.00382$ $0.00931\pm 0.00837$
20000 4 11 $4.155\pm 0.00758$ $0.00887\pm 0.00792$
200000 4 11 $41.869\pm 0.0203$ $0.00880\pm 0.00777$
2000000 4 11 $381.313\pm 0.199$ $0.0114\pm 0.00803$
Table A.3: Average wall-clock runtime of Algorithm 1 for different values of $N$ (the number of data points), $K$ (the covering set size), and $T$ (the number of objectives). The “Runtime" column gives the average runtime of Algorithm 1 with parallel computation of the of the marginal coverage of adding the $N$ data points (parallel computation of lines 5-6). The “Runtime No Parallelization" column gives the average runtime of Algorithm 1 without this parallelization of the inner for loop. Average run times are computed by running Algorithm 1 on a NVIDIA RTX A5000 GPU and averaging over $10$ runs. Standard errors over the $10$ runs are also provided.
Efficiency of Algorithm 1
Algorithm 1 could also be made more efficient by pruning the points in $D_{s}$ with zero marginal coverage improvement after each outer-loop, since these points will continue to have zero marginal coverage improvement on subsequent loops. In future work, we plan to explore this and any other tricks that might allow us to further improve the efficiency of Algorithm 1.
A.3 Approximation strategy for batch acquisition with Expected Coverage Improvement (ECI)
In batch acquisition, we select a batch of $q>1$ candidates for evaluation. In Equation 6, we define q-ECI, a natural extension of the ECI acquisition function defined in Equation 4 to the batch acquisition setting. q-ECI gives the expected improvement in the coverage score after simultaneously observing the batch of $q$ points $\mathbf{X}=\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{q}\right\}$. When using batch acquisition, we aim to select a batch of $q$ points that maximize q-ECI.
We will first discuss how one would estimate q-ECI using a Monte Carlo (MC) approximation. To select a batch candidates $\hat{\mathbf{X}}$, we sample $m$ batches of $q$ points $B=\left\{\mathbf{B}_{1},\mathbf{B}_{2},...,\mathbf{B}_{m}\right\}$. Here $\mathbf{B}_{j}$ is a batch of $q$ sampled points $\mathbf{B}_{j}=\left\{\mathbf{b}_{j1},\ldots,\mathbf{b}_{jq}\right\}$. For each batch $\mathbf{B}_{j}$, we sample a realization $\hat{\mathbf{Y}}_{j}=\left\{\hat{\mathbf{y}}_{j1},\ldots,\hat{\mathbf{y}}_{jq}\right\}$ from the GP surrogate model posterior. We leverage these samples to compute an MC approximation to the q-ECI of each $\mathbf{B}_{j}$:
$q\text{-CI}(\mathbf{B}_{j})$ $=\max(0,\max_{r=1,\dots,q}c(S^{*}_{D_{s}\cup\left\{(\mathbf{b}_{jr},\hat{\mathbf{y}}_{jr})\right\}})-c(S^{*}_{D_{s}}))$ (7)
[…]
Instead of sampling $m$ batches of candidates, we sample $m$ individual data points $P=\left\{\mathbf{p}_{1},\mathbf{p}_{2},...,\mathbf{p}_{m}\right\}$. For each sampled point $\mathbf{p}_{j}$, we sample a realization $\hat{\mathbf{y}}_{j}=(\hat{f}_{1}(\mathbf{p}_{j}),\ldots,\hat{f}_{T}(\mathbf{p}_{j}))$. As in Section 3.1.1, we use the sampled realizations $\hat{\mathbf{y}}_{j}$ to compute an approximate coverage improvement $CI(\mathbf{p}_{j})$ as defined in Equation 5 for each point $\mathbf{p}_{j}$. To obtain a batch of $q$ candidates, we then greedily select the $q$ points $\mathbf{p}_{j}\in P$ with the $q$ largest expected coverage improvements. Note that this strategy does not involve sequential optimization of the $q$ points, as the batch of $q$ points is selected simultaneously as the points with the top-$q$ expected coverage improvements.
A.4 Limitations and future work
A primary limitation of MOCOBO is that the choice of $K$ (the covering set size) may not always be straightforward. For example, as we mention in Section 4, it may require domain knowledge to choose $K<T$ that is large enough that achieving good coverage is possible despite multiple conflicting objectives. In many practical applications, we prefer the smallest possible set size $K$ such that good coverage can still be achieved. In future work, we plan to explore methods for simultaneously optimizing both the solutions in the covering set, and the size of the covering set, balancing the trade-off between minimizing the number of solutions needed and maximizing overall coverage.
A.5 Broader impacts
This research includes applications in molecule and peptide design. While AI-driven biological design holds great promise for benefiting society, it is crucial to acknowledge its dual-use potential. Specifically, AI techniques designed for drug discovery could be misused to create harmful biological agents [65].
Our goal is to accelerate drug development by identifying promising candidates, but it is imperative that experts maintain oversight, that all potential therapeutics undergo thorough testing and clinical trials, and that strict regulatory frameworks governing drug development and approval are followed.
A.6 Compute resources
In this section, we provide all details about compute resources used to produce all results in this paper.
Table A.4: Setup of internal cluster used to run experiments.
Type Specifications
System Topology 20 nodes with 2 sockets each with 24 logical threads (total 48 threads)
Processor 1 Intel Xeon Silver 4310, 2.1 GHz (maximum 3.3 GHz) per socket
Cache 1.1 MiB L1, 30 MiB L2, and 36 MiB L3
Memory 250 GiB RAM
Accelerator 1 NVIDIA RTX A5000 per node, 2 GHZ, 24GB RAM
Compute specifications (type and memory).
We use GPU works to run all experiments and produce all empirical results provided in this paper. A single GPU was used per run of each method compared on each task. Each each run uses approximately 12-18 GB of the GPU memory. Most experiments were executed on our internal cluster of NVIDIA RTX A5000 GPUs (see internal cluster compute details in Table A.4). We also used cloud compute resources for two weeks to complete additional replications of some experiments. We used a total of eight RTX 4090 GPU workers from runpod.io, each with approximately 24 GB of GPU memory.
Execution time.
For the relatively inexpensive rover task, each optimization run takes approximately 1 day of execution time. For all other tasks considered, each optimization run takes approximately 3 days of execution time. To create all coverage optimization plots, we ran all methods compared $20$ times each. Completing all of the runs needed to produce all of the results in this paper required roughly $64000$ total GPU hours.
Compute resources for preliminary experiments.
Preliminary experiments refer to the initial experiments for e.g. method development that are not included as results in the paper. All preliminary experiments were done our internal cluster of NVIDIA RTX A5000 GPUs (see internal cluster compute details in Table A.4). We spent approximately $2000$ hours of GPU time on preliminary experiments.
A.7 Additional implementation details
In this section, we provide additional implementation details for MOCOBO. After the paper is accepted, we will also refer readers to the MOCOBO codebase for the full-extent of implementation details and experimental setup needed to reproduce the results provided https://github.com/placeholder.
A.7.1 Trust region hyperparameters
For all trust region methods, the trust region hyperparameters are set to the TuRBO defaults used by Eriksson et al. [17].
A.7.2 Surrogate model
Since the tasks considered in this paper are challenging high-dimensional tasks, requiring a large number of function evaluations, we use approximate GP surrogate models. In particular, we use a PPGPR [28] surrogate models with a constant mean, standard RBF kernel, and $1024$ inducing points. Additionally, we use a deep kernel (several fully connected layers between the search space and the GP kernel) [70]. We use two fully connected layers with $D$ nodes each, where $D$ is the dimensionality of the search space.
We use the same PPGPR model(s) with the same configuration for MOCOBO, TuRBO, LOL-BO, and ROBOT. For MOCOBO, to model the $T$-dimensional output space, we use $T$ PPGPR models, one to approximate each objective $f_{1},\ldots,f_{T}$. To allow information sharing between the models, we use a shared deep kernel (the $T$ PPGPR models share share the same two-layer deep kernel) [29, 49].
Unlike the other methods compared, MORBO was designed for use with an exact GP model rather than an approximate GP surrogate model. For fair comparison, we therefore run MORBO with an exact GP using all default hyperparameters and the official codebase provided by Daulton et al. [11].
We train the PPGPR surrogate model(s) on data collected during optimization using the Adam optimizer with a learning rate of $0.001$ and a mini-batch size of $256$. On each step of optimization, we update the model on collected data until we stop making progress (loss stops decreasing for $3$ consecutive epochs), or exceed $30$ epochs. Since we collect a large amount of data for each optimization run (i.e. as many as $62e$ 6 data points in a single for the “template constrained" peptide design task), we avoid updating the model on all data collected at once. On each step of optimization, we update the current surrogate model only on a subset of $1000$ of the collected data points. This subset is constructed of the data that has obtained the highest objective values so far, along with the most recent batch of data collected. By always updating on the most recent batch of data collected, we ensure that the surrogate model is conditioned on every data point collected at some point during the optimization run.
A.7.3 Initialization data
In this section, we provide details regarding the data used to initialize all optimization runs for all tasks in Section 4.
To initialize optimization for the molecule design task, we take a random subset of $10000$ molecules from the standardized unlabeled dataset of 1.27M molecules from the Guacamol benchmark software [6]. We generate labels for these $10,000$ molecules once, and then use the labeled data to initialize optimization for all methods compared.
To initialize optimization for the peptide design tasks, we generate a a set of $20,000$ peptide sequences by making random edits (insertions, deletions, and mutations) to the $10$ template peptide sequences in Table A.8. We generate labels for these $20,000$ peptides once, and then use the labeled data to initialize optimization for all methods compared.
For all other tasks, we initialize optimization with $2,000$ points sampled uniformly at random from the search space.
A.7.4 Diversity constraints and associated hyperparameters for ROBOT baseline
For a single objective, ROBOT seeks a diverse set of $M$ solutions, requiring that the set of solutions have a minimum pairwise diversity $\tau$ according to the user specified diversity function $\delta{}$. Since Maus et al. [45] also consider rover and molecule design tasks, we use the same diversity function $\delta{}$ and diversity threshold $\tau$ used by Maus et al. [45] for these two tasks. For the peptide design tasks, we define $\delta{}$ to be the edit distance between peptide sequences, and use a diversity threshold of $\tau=3$ edits. For the image optimization task, since there is no obvious semantically meaningful measure of diversity between two sets of input parameters, we define $\delta{}$ to be the euclidean distance between solutions, and use $\tau=1.45$, the approximate average euclidean distance between a randomly selected pair of points in the search space.
A.8 Additional task details
In this section we provide additional details for the chosen set of tasks we provide results for in Section 4.
A.8.1 HDR image tone mapping
Objective ID Pyiqa Metric ID Reference
1 nima 59
2 nima-vgg16-ava 59, 46
3 topiq-iaa-res50 10
4 laion-aes 53
5 hyperiqa 58
6 tres 22
7 liqe 72
Table A.5: pyiqa metric ID strings used to identify the $T=7$ target image quality metrics used for image tone mapping. Each metric is a no-reference image aesthetic (IAA) or quality (IQA) assessment metric obtained from the pyiqa library [9].
In Table A.5, we list the names of the $7$ image quality metrics used for the image tone mapping tasks described in Section 4.1.
Target metrics.
The target image quality (IQA) and aesthetic (IAA) metrics are organized in Table A.5, where all except nima are IAA metrics, while nima is an IQA metric. For more detailed information about each metric and their corresponding datasets, please refer to their original references.
Benchmark images.
We use two benchmark images. The first is the “Stanford Memorial Church" image obtained from https://www.pauldebevec.com/Research/HDR/ by courtesy of Paul E. Debevec [12]. The second is the “desk lamp" image obtained from https://cadik.posvete.cz/tmo/ by courtesy of Martin Čadík [7]. Because commonly used metrics are only correlated with subjective image quality, prior work on tuning parameters for these and related benchmarks has been done by trial and error and human-in-the-loop type schemes [41], including preferential BO-based approaches [35, 36].
Imaging pipeline.
We consider a tone mapping pipeline consisting of a multi-layer detail decomposition [62, 14, 38] using the guided filter by He et al. [25] (3 detail layers and 1 base layer), followed by gamma correction [51, Section 2.9], resulting in a 13-dimensional optimization problem. The complete image processing pipeline is very similar to the classic approach proposed by Tumblin and Turk [62], where the main difference is that, similarly to Farbman et al. [18], we replace the diffusion smoothing filter with a more recent edge-preserving detail smoothing filter, the guided filter, by He et al. [25]. First, given an HDR image in the RGB color space $I=(I_{\mathrm{r}},I_{\mathrm{r}},I_{\mathrm{b}}))$, where $I_{\mathrm{r}}$ is the red channel, $I_{\mathrm{r}}$ is the blue channel, and $I_{\mathrm{g}}$ is the green channel, we compute the luminance according to
$L\triangleq 0.2989I_{r}+0.587I_{g}+0.114I_{b}\,.$
(The constants were taken from the code of Farbman et al. 18.) The luminance image $L$ is then logarithmically compressed and then decomposed into three detail layers and one base layer by applying the guided filter three times, each with a different radius parameter $r_{i}$ and smoothing parameter $\epsilon_{i}$ for $i=1,\ldots,3$. Then, we amplify or attenuate each channel with a corresponding gain coefficient $g_{\mathrm{detail},i}$ $i=1,\ldots,3$ for the detail and $g_{\mathrm{base}}$ for the base layers. The image is then reconstructed by adding all the layers, including the colors. Following Tumblin and Turk [62], we also apply a gain, $g_{\mathrm{color}}$, to the color channels. Finally, the resulting image is applied an overall gain $g_{\mathrm{out}}$ and then gamma-corrected [51, Section 2.9] with an exponent of $1/\gamma$. The parameters for this pipeline are organized in Table A.6. The implementation uses OpenCV [5], in particular, the guided filter implementation in the extended image processing (ximgproc) submodule.
Parameter Description Domain
$r_{i}$ radius of the guided filter for generating the $i$th detailed layer $\{3,\ldots,32\}$
$\epsilon_{i}$ $\epsilon$ of the guided filter for generating the $i$th detailed layer $[0.01,10]$
$g_{\mathrm{detail},i}$ Gain of the $i$th detail layer $[0,1.5]$
$g_{\mathrm{base}}$ Gain of the $i$th detail layer $[0,1]$
$g_{\mathrm{color}}$ Gain of the color layer $[0.5,1.5]$
$g_{\mathrm{out}}$ Gain of tone-mapped output $[0.2,2.0]$
$\gamma$ Gamma correction inverse exponent $[1,5]$
Table A.6: pyiqa metric ID strings used to identify the $T=7$ target image quality metrics used for the image tone mapping task.
Optimization problem setup.
For optimization, we map the parameters
$\mathbf{x}=\left(r_{1},\,r_{2},\,r_{3},\,\epsilon_{1},\,\epsilon_{2},\,\epsilon_{3},\,g_{\mathrm{detail},1},\,g_{\mathrm{detail},2},\,g_{\mathrm{detail},3},\,g_{\mathrm{base}},\,g_{\mathrm{color}},\,g_{\mathrm{out}},\,\gamma\right)$
to the unit hypercube ${[0,1]}^{13}$. In particular, the mapped values on the unit interval $[0,1]$ linearly interpolate the domain of each parameter shown in Table A.6. For the radius parameters $r_{1},r_{2},r_{3}$, which are categorical, we naively quantize the domain by rounding the output of the interpolation to the nearest integer.
A.8.2 Peptide design
Objective ID Target Pathogenic Bacteria
B1 A. baumannii ATCC 19606
B2 E. coli ATCC 11775
B3 E. coli AIC221
B4 E. coli AIC222-CRE
B5 K. pneumoniae ATCC 13883
B6 P. aeruginosa PAO1
B7 P. aeruginosa PA14
B8 S. aureus ATCC 12600
B9 S. aureus ATCC BAA-1556-MRSA
B10 E. faecalis ATCC 700802-VRE
B11 E. faecium ATCC 700221-VRE
Table A.7: Names of the $T=11$ target bacteria used for the peptide design task. The first seven bacteria are Gram negative (IDs B1-B7) and the last four (IDs B8-B11) are Gram positive.
Template Amino Acid Sequences
RACLHARSIARLHKRWRPVHQGLGLK
KTLKIIRLLF
KRKRGLKLATALSLNNKF
KIYKKLSTPPFTLNIRTLPKVKFPK
RMARNLVRYVQGLKKKKVI
RNLVRYVQGLKKKKVIVIPVGIGPHANIK
CVLLFSQLPAVKARGTKHRIKWNRK
GHLLIHLIGKATLAL
RQKNHGIHFRVLAKALR
HWITINTIKLSISLKI
Table A.8: Template amino acid sequences used for the “template constrained" peptide design task.
Table A.7 specifies the $T=11$ target bacteria used for the peptide design task from Section 4. The first seven bacteria are Gram negative bacteria (Objective IDs B1-B7) and the last four (Objective IDs B8-B11) are Gram positive. Table A.8 gives the $10$ template amino acid sequences used for the “template constrained" variation of the peptide design task from Section 4.
A.9 Proof that finding $S^{*}_{D_{s}}$ is NP-hard
In this section, we prove Lemma 3.1, that finding $S^{*}_{D_{s}}$ Is NP-Hard.
Proof.
We prove Lemma 3.1 this by reduction from the well-known Maximum Coverage Problem (MCP), which is NP-Hard.
Definition A.1 (Maximum Coverage Problem (MCP)).
In the Maximum Coverage Problem, we are given:
  • A universe $U=\{e_{1},e_{2},\ldots,e_{m}\}$ of $m$ elements.
  • A collection of $n$ subsets $\mathcal{S}=\{A_{1},A_{2},\ldots,A_{n}\}$, where $A_{i}\subseteq U$.
  • An integer $K$, the number of subsets we can select.
The objective is to find a collection of $K$ subsets $\mathcal{S}^{\prime}\subseteq\mathcal{S}$ such that the total number of elements covered, $\bigcup_{A\in\mathcal{S}^{\prime}}A$, is maximized.
Proposition A.2 (MCP is NP-Hard).
The Maximum Coverage Problem is NP-Hard.
Reduction from MCP to Finding $S^{*}_{D_{s}}$:
We reduce an instance of MCP to the problem of finding $S^{*}_{D_{s}}$ as follows:
  1. 1.
    Let the universe $U=\{e_{1},e_{2},\ldots,e_{m}\}$ correspond to the objectives $\{1,2,\ldots,T\}$ in the optimal covering set problem, i.e., set $T=m$.
  2. 2.
    Let each subset $A_{i}\in\mathcal{S}$ correspond to a point $\mathbf{x}_{i}\in D_{s}$.
  3. 3.
    Define $f_{t}(\mathbf{x}_{i})=1$ if subset $A_{i}$ covers element $e_{t}$, and $f_{t}(\mathbf{x}_{i})=0$ otherwise.
  4. 4.
    The goal of the Maximum Coverage Problem (selecting $K$ subsets to maximize coverage) now translates to selecting $K$ points $\mathbf{x}_{i}\in D_{s}$ to maximize the coverage score:
    $c(S)=\sum_{t=1}^{T}\max_{\mathbf{x}\in S}f_{t}(\mathbf{x}).$
Correctness of the Reduction:
The reduction ensures that:
  • Each subset $A_{i}\in\mathcal{S}$ corresponds to a point $\mathbf{x}_{i}\in D_{s}$.
  • If subset $A_{i}$ covers element $e_{t}$, then $f_{t}(\mathbf{x}_{i})=1$; otherwise, $f_{t}(\mathbf{x}_{i})=0$.
  • Selecting $K$ subsets in MCP to maximize coverage corresponds exactly to selecting $K$ points $\mathbf{x}_{i}$ in $D_{s}$ to maximize the coverage score $c(S)$, which sums the contributions of all objectives (covered elements).
Thus, solving the optimal covering set problem is equivalent to solving MCP.
Implications:
Since MCP is NP-Hard (Proposition A.2), and we have reduced MCP to the problem of finding $S^{*}_{D_{s}}$ in polynomial time, it follows that finding $S^{*}_{D_{s}}$ is also NP-Hard.
Proof significance
We do not claim the above proof of Lemma 3.1 as a novel contribution of this work, as it follows straightforwardly from the fact that the well-known Maximum Coverage Problem is NP-hard. Note that proving Lemma 3.1 would also follow straightforwardly from the work of Ding et al. [13] who demonstrate that the coverage optimization problem generalizes k-means clustering. Rather than providing additional novel contribution, the above proof of Lemma 3.1 serves to justify our use of a greedy approximation algorithm (Algorithm 1) to approximate $S^{*}_{D_{s}}$ on each iteration of MOCOBO.
A.10 Approximation proof for Algorithm 1
In this section, we prove Theorem 3.2, that Algorithm 1 is a $(1-\frac{1}{e})$-Approximation.
Definition A.3 (Coverage Score).
The coverage score of a set $S\subseteq D_{s}$, denoted $c(S)$, is defined as:
$c(S)=\sum_{t=1}^{T}\max_{\mathbf{x}\in S}f_{t}(\mathbf{x}),$
where $f_{t}(\mathbf{x})$ is the observed value of objective $t$ at point $\mathbf{x}$.
Definition A.4 (Optimal Covering Set).
Let $S^{*}_{D_{s}}\subseteq D_{s}$ denote the optimal covering set of size $K$:
$S^{*}_{D_{s}}=\operatorname*{arg\,max}_{S\subseteq D_{s},|S|=K}c(S).$
Its coverage score is given by $c(S^{*}_{D_{s}})$.
Definition A.5 (Contribution of Objective Function $f_{t}$ to Coverage Score).
Given a covering set $S\subseteq D_{s}$, we denote the contribution of objective $f_{t}\in\{f_{1},f_{2},\ldots,f_{T}\}$ to the overall coverage score $c(S)$ as $g_{t}(S)$ where
$g_{t}(S)=\max_{\mathbf{x}\in S}f_{t}(\mathbf{x})$
It follows that from Definition A.5 that:
$c(S)=\sum_{t=1}^{T}g_{t}(S)$
Lemma A.6 (Monotonicity).
The coverage score $c(S)$ as defined in Definition A.3 is monotone, i.e., for any $S\subseteq S^{\prime}\subseteq D_{s}$, we have:
$c(S)\leq c(S^{\prime}).$
Proof.
Adding more points to a set can only increase or maintain the maximum values of $f_{t}$ for each objective $t$, since for each objective:
$g_{t}(S^{\prime})=\max\left(g_{t}(S),g_{t}(S^{\prime}\setminus S)\right)\geq g_{t}(S).$
Hence, $c(S)$ is monotone. ∎
Lemma A.7 (Submodularity).
The coverage score $c(S)$ as defined in Definition A.3 is submodular, i.e., for any $S\subseteq S^{\prime}\subseteq D_{s}$ and any $\mathbf{x}^{*}\in D_{s}\setminus S^{\prime}$, we have:
$c(S\cup\{\mathbf{x}^{*}\})-c(S)\geq c(S^{\prime}\cup\{\mathbf{x}^{*}\})-c(S^{\prime}).$
Proof.
The marginal improvement of adding $\mathbf{x}^{*}$ to $S$ is:
$c(S\cup\{x^{*}\})-c(S)$ $=\sum_{t=1}^{T}\left(g_{t}(S\cup\{x^{*}\})-g_{t}(S)\right)$
$=\sum_{t=1}^{T}\max\left[f_{t}(x^{*}),g_{t}(S)\right]-g_{t}(S)$
$=\sum_{t=1}^{T}\begin{cases}f_{t}(\mathbf{x}^{*})-g_{t}(S)&f_{t}(\mathbf{x}^{*})>g_{t}(S)\\ 0&\textrm{otherwise}\end{cases}$
and likewise for $S^{\prime}$ this is:
$c(S^{\prime}\cup\{x^{*}\})-c(S^{\prime})=\sum_{t=1}^{T}\begin{cases}f_{t}(\mathbf{x}^{*})-g_{t}(S^{\prime})&f_{t}(\mathbf{x}^{*})>g_{t}(S^{\prime})\\ 0&\textrm{otherwise}\end{cases}$
Now, noting that for each term
$g_{t}(S^{\prime})=\max\left[g_{t}(S),g_{t}(S^{\prime}\setminus S)\right]\geq g_{t}(S),$
we know then that $f_{t}(\mathbf{x}^{*})-g_{t}(S^{\prime})\leq f_{t}(\mathbf{x}^{*})-g_{t}(S)$, and further that $f_{t}(\mathbf{x}^{*})>g_{t}(S^{\prime})$ implies that $f_{t}(\mathbf{x}^{*})>g_{t}(S)$. Taken together, these imply that:
$\sum_{t=1}^{T}\begin{cases}f_{t}(\mathbf{x}^{*})-g_{t}(S^{\prime})&f_{t}(\mathbf{x}^{*})>g_{t}(S^{\prime})\\ 0&\textrm{otherwise}\end{cases}\leq\sum_{t=1}^{T}\begin{cases}f_{t}(\mathbf{x}^{*})-g_{t}(S)&f_{t}(\mathbf{x}^{*})>g_{t}(S)\\ 0&\textrm{otherwise}\end{cases}$
and therefore:
$c(S\cup\{\mathbf{x}^{*}\})-c(S)\geq c(S^{\prime}\cup\{\mathbf{x}^{*}\})-c(S^{\prime}).$
Lemma A.8 (Greedy Achieves a $(1-\frac{1}{e})$-Approximation for Monotone Submodular Functions).
For any monotone submodular function, the greedy submodular optimization strategy provides a $(1-\frac{1}{e})$-approximation.
Proof.
This is a well know result about monotone submodular functions, show for example by Nemhauser et al. [48]. ∎
Proof.
Proof of Theorem 3.2 From Lemma A.6 and Lemma A.6, we know that $c(S)$ is monotone submodular.
Since $c(S)$ is monotone submodular, and Algorithm 1 approximates $A^{*}_{D_{s}}$ using greedy submodular optimization, it follows from Lemma A.8 that Algorithm 1 achieves:
$c(A^{*}_{D_{s}})\geq\left(1-\frac{1}{e}\right)c(S^{*}_{D_{s}})$
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: All stated claims are backed-up with results in Section 4 and the stated focus/scope of the paper accurately reflects what is discussed throughout the rest of the paper.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: See Section A.4 for discussion of limitations.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: All theoretical claims can be found in Section 3.1.2. All sets of assumptions and full proofs of each claim are provided in Section 3.1.2, Section A.9, and Section A.10. As far as we are aware, each proof is complete and correct.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We provide detailed explanation of how our method works in Section 3 and all additional required details to reproduce results in Section 4 and Section A.7. Additionally, once the paper is accepted, we will replace the placeholder GitHub links with a link to a public GitHub repository containing the source code used to produce the results provided in Section 4. This source code will allow any reader to run our code to reproduce results on all tasks in the paper. Additionally, the README in the repository provides detailed instructions to make setting up the proper environment and running the code easy for users.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: Once the paper is accepted, we will replace the placeholder GitHub links in the main text with a link to a public GitHub repository containing the source code used to produce the results provided in this paper. This GitHub repository contains organized code that will allow any reader to run MOCOBO on all tasks to reproduce the results provided in this paper. Additionally, the README in the repository provides detailed instructions to make setting up the proper environment and running the code easy for users.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: All chosen hyper-parameters and implementation details are stated in section 4 and Section A.7.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: On all plots, we plot the mean taken over $20$ random runs and include error bars to show the standard error over the runs.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: All compute details are provided in Section A.6.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We have made sure to adhere to the NeurIPS Code of Ethics in all aspects of our research.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: See discussion of broader societal impacts in Section A.5.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The paper does not release new data or models with potential societal concerns.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: Creators of all assets used to produce all results in this paper are cited in Section 4. All assets used are open source software or models.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: This paper does not introduce any new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The work does not involve any human participants.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The work does not involve any living participants.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: This work does not use LLMs in any way.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.